# Do Robots Really Need Anthropomorphic Hands?

### v1

@@ 5.1.1 Do we need an opposable thumb?
The thumb is the most independently controlled of the human fingers [33]. An opposable thumb improves the prehensile function of a hand compared to parallel fingers, as it allows us to exert forces on an object from opposing directions. However, a robotic hand does not necessarily need a thumb that reflects the function and shape of the human thumb. Parallel jaw grippers, featuring two opposing fingers, are capable of producing stable grasps. It is not even necessary to have more than one finger to apply forces from opposing directions, e.g., it is possible to hold an object between one finger and a palm, or an octopus arm can wrap around an object. The BarrettHand has three fingers, and two of them can rotate 180 degrees around their base. This capability enables configurations in which one finger is opposing the others, in which all three fingers are equally distributed around the object, or in which all fingers are parallel. [8] point out that the human thumb creates an asymmetry and constrains the orientation of the hand. They entertain the idea of designing hands with more than one opposing finger to enable more dexterous in-hand manipulation, which would usually require two hands. There are examples of such hands with two pairs of opposing fingers [66, 55].
@@
Comparing independent finger movements of humans versus monkeys, [30] found that individuation is higher in human subjects. Specifically, they found that the thumb and index fingers are the most independently moving fingers, while the ring and middle fingers are most coupled with the movement of other fingers. This observation coincides with the fact that the thumb, index finger, and little finger have more muscles for individual actuation of these fingers. The humans’ greater independence of finger movements in comparison to macaques, for instance, arises from anatomical differences such as splitting of a separate muscle belly and tendons, and a loss of tendons in multitendoned muscles [30], but also from finer representation in the primary motor cortex. Nevertheless, humans still often do not control each finger independently. [30] found that the angular motion mostly happens in the middle joint of each finger (usually PIP joint, MCP for the thumb). This observation indicates that a reduced set of actively controlled flexion joints might be sufficient.
By constraining individual fingers of human subjects during manipulation experiments, [63] explored the relevance of independent fingers in activities of daily living. Their results suggest that the precision grasps of the human hand with small contact areas rely on independent abduction/adduction of the fingers more than on independent flexion/extension, as this allows for precise opposition of the thumb.
@@ 5.1.4 Do we need a soft and deformable surface, particularly at the large surface of the flexible palm?
[1] measure the distribution of forces for different grasp types. Although the only force sensors on the palm that they use are below the little finger, their heatmaps suggest that the palm is important for human grasps. The benefit of flexible and soft hands is twofold: (1) contact with the environment does not damage the hardware, and (2) support of prehension by adaptation to the shape of objects. Humans do not avoid contact with the environment during manipulation; they actively use it and exploit environmental constraints. For example, when grasping a flat object from a table, they often slide it to the edge to grasp it more easily [20].
@@ 6 Conclusion
In the previous section, we provided recommendations for designing robotic hands. For prosthetics, emulating the mechanical characteristics of the human hand is a sensible goal. For robots, however, this is less applicable. With intelligent control strategies, even simple two- or three-fingered grippers can perform tasks effectively, for example, by leveraging the environment to counteract forces from other directions. We can even go beyond human manipulation abilities. More complex, non-anthropomorphic hands can solve complex manipulation problems more easily or multiple problems simultaneously because they can directly control many DoFs of one or multiple objects without repositioning the hand. They could also solve various manipulation tasks simultaneously. However, how this can be done remains an open question.
@@
However, are we, as roboticists, actually looking at the right defining characteristics of human hands? While we can roughly match the degrees of freedom and number of fingers of a human hand, we are not yet able to match the force-to-weight ratio in dexterous hands. This fact is well-known, but this may not be the most interesting aspect of a human hand. We assume that learning and control are a large part of what is currently missing for human-like dexterity. However, learning is a form of iteratively improving stochastic control that is only possible with robust hardware. The human hand is soft, deformable, and flexible, with a larger passive range of motion, which makes it particularly robust and suited for approximate control and manipulation under environmental uncertainty. We argue that we optimized hand designs for the wrong characteristics of human hands in the past. We can perform complex skills with three or four fingers and fewer degrees of freedom. Robustness to environmental conditions should be the focus of hand design.
@@
The synergetic combination of mechanical aspects, perception, and control might be more important for a robotic hand to match and potentially surpass human dexterity than anthropomorphic kinematics alone. Teleoperation demonstrates that the most limiting factor currently is the cognitive ability of robotic systems, as surgeons can perform complex procedures using simple grippers and tools without haptic feedback. However, haptic feedback, even in its current state of development, has a significant overall effect on measures of applied forces, completion time, accuracy, and success rates [5].
@@
Our systematic review reveals a key shortcoming: most hands are tested within only a few categories of Dollar’s taxonomy [21], and even within these categories, there is limited variety. This lack of comprehensive evaluation is especially evident in cases of complex hands, which are seldom tested for in-hand manipulation. To address this, the field needs robust evaluation protocols that can assess complete mechanism-sensor-software systems across all skill categories and complexity levels. Such protocols must demonstrate a hand’s mechanical and functional capabilities, especially by requiring the intelligent use of advanced sensors. For example, to truly evaluate the benefits of tactile sensors, protocols should include complex manipulation tasks that demand fine motor control. Establishing these benchmarks is crucial for making meaningful comparisons of hand designs and for guiding progress in robotic hand development.
[…]
Supplementary information An overview of the analyzed papers is available at https://figshare.com/s/f4fdd9aca1c134f010dd.
Declarations
[…]
Open Access This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution, and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons license, and indicate if changes were made. The images or other third-party material in this article are included in the article’s Creative Commons license, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons license and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this license, visit http://creativecommons.org/licenses/by/4.0/.
[…]
@@ Appendix A Radar Diagram of Skills
  • Sign language: For ”I love you” in American sign language, five fingers are usually required, these have to provide flexion/extension DOFs, and the thumb has to be in reposition.

### v2

@@ 5.1.1 Do we need an opposable thumb?
The thumb is the most independently controlled of the human fingers [13]. An opposable thumb improves the prehensile function of a hand compared to parallel fingers, as it allows us to exert forces on an object from opposing directions. However, a robotic hand does not necessarily need a thumb that reflects the function and shape of the human thumb. Parallel jaw grippers, featuring two opposing fingers, are capable of producing stable grasps. It is not even necessary to have more than one finger to apply forces from opposing directions, e.g., it is possible to hold an object between one finger and a palm, or an octopus arm can wrap around an object. The BarrettHand has three fingers, and two of them can rotate 180 degrees around their base. This capability enables configurations in which one finger is opposing the others, in which all three fingers are equally distributed around the object, or in which all fingers are parallel. Billand and Kragic [4] point out that the human thumb creates an asymmetry and constrains the orientation of the hand. They entertain the idea of designing hands with more than one opposing finger to enable more dexterous in-hand manipulation, which would usually require two hands. There are examples of such hands with two pairs of opposing fingers [110, 111].
@@
Comparing independent finger movements of humans versus monkeys, Häger-Ross and Schieber [125] found that individuation is higher in human subjects. Specifically, they found that the thumb and index fingers are the most independently moving fingers, while the ring and middle fingers are most coupled with the movement of other fingers. This observation coincides with the fact that the thumb, index finger, and little finger have more muscles for individual actuation of these fingers. The humans’ greater independence of finger movements in comparison to macaques, for instance, arises from anatomical differences such as splitting of a separate muscle belly and tendons, and a loss of tendons in multitendoned muscles [125], but also from finer representation in the primary motor cortex. Nevertheless, humans still often do not control each finger independently. Häger-Ross and Schieber [125] found that the angular motion mostly happens in the middle joint of each finger (usually PIP joint, MCP for the thumb). This observation indicates that a reduced set of actively controlled flexion joints might be sufficient.
By constraining individual fingers of human subjects during manipulation experiments, Montagnani et al. [114] explored the relevance of independent fingers in activities of daily living. Their results suggest that the precision grasps of the human hand with small contact areas rely on independent abduction/adduction of the fingers more than on independent flexion/extension, as this allows for precise opposition of the thumb.
@@ 5.1.4 Do we need a soft and deformable surface, particularly at the large surface of the flexible palm?
Abbasi et al. [117] measure the distribution of forces for different grasp types. Although the only force sensors on the palm that they use are below the little finger, their heatmaps suggest that the palm is important for human grasps. The benefit of flexible and soft hands is twofold: (1) contact with the environment does not damage the hardware, and (2) support of prehension by adaptation to the shape of objects. Humans do not avoid contact with the environment during manipulation; they actively use it and exploit environmental constraints. For example, when grasping a flat object from a table, they often slide it to the edge to grasp it more easily [113].
@@ 6 Conclusion
In the previous section, we provided recommendations for designing robotic hands. For prosthetics, emulating the mechanical characteristics of the human hand is a sensible goal. For robots, however, this is less applicable and may even be counterproductive in human-robot interaction because of the uncanny valley [134]. With intelligent control strategies, even simple two- or three-fingered grippers can perform tasks effectively, for example, by leveraging the environment to counteract forces from other directions. Instead of increasing the fingers and individually controllable flexion/extension joints, we recommend to mimic the flexibility of the human arm with seven DoF, which allows to avoid obstacle in the workspace or torque limits, or to add abduction/adduction joints for of each finger, which allows exact opposition and, hence, precision grasps with small contact areas.
We can even go beyond human manipulation abilities. More complex, non-anthropomorphic hands can solve complex manipulation problems more easily or multiple problems simultaneously because they can directly control many DoFs of one or multiple objects without repositioning the hand. They could also solve various manipulation tasks simultaneously. However, how this can be done remains an open question.
@@
Are we, as roboticists, actually looking at the right defining characteristics of human hands? While we can roughly match the degrees of freedom and number of fingers of a human hand, we are not yet able to match the force-to-weight ratio in dexterous hands. This fact is well known, but it may not be the most interesting aspect of a human hand. We assume that learning and control are a large part of what is currently missing for human-like dexterity. However, learning is a form of iteratively improving stochastic control that is only possible with robust hardware. The human hand is soft, deformable, and flexible, with a larger passive range of motion, making it particularly robust and well-suited for approximate control and manipulation under environmental uncertainty. We may have optimized hand designs for the wrong characteristics of human hands in the past, because, as discussed in this article, we can perform many skills with less fingers or fewer degrees of freedom than the human hand. Robustness to potential damages, for instance, overstretching of joints or collisions with the environment, should be the focus of hand design to enable interactive sensorimotor learning, i.e., trial and error, for a wide range of skills.
@@
The synergetic combination of mechanical aspects, perception, and control might be more important for a robotic hand to match and potentially surpass human dexterity than anthropomorphic mechanisms alone, suggesting that form-based biomimicry is inferior to function-based biomimicry. Teleoperation demonstrates that the most limiting factor currently is the cognitive ability of robotic systems, as surgeons can perform complex procedures using simple grippers and tools without haptic feedback. However, haptic feedback, even in its current state of development, has a significant overall effect on measures of applied forces, completion time, accuracy, and success rates [135].
@@
Our systematic review reveals a key shortcoming of the current state of the art: most hands are tested within only a few categories of Dollar’s taxonomy [108], and even within these categories, there is limited variety. This lack of comprehensive evaluation is especially evident in cases of complex hands, which are seldom tested for in-hand manipulation. To address this, the field needs robust evaluation protocols that can assess complete mechanism-sensor-software systems across all skill categories and complexity levels. Such protocols must demonstrate a hand’s mechanical and functional capabilities, especially by requiring the intelligent use of advanced sensors. For example, to truly evaluate the benefits of tactile sensors, protocols should include complex manipulation tasks that demand fine motor control. Establishing these benchmarks is crucial for making meaningful comparisons of hand designs and for guiding progress in robotic hand development.
Acknowledgments
This work was supported by the European Commission under the Horizon 2020 framework program for Research and Innovation via the APRIL project (project number: 870142) and by the Vibro-Sense Project (project number: 03DPS1242A) funded by the Bundesministerium Forschung, Technologie und Raumfahrt (BMFTR) under the DATIpilot program. Open Access funding provided by the Projekt DEAL (Open access agreement for Germany).
@@ Appendix A Radar Diagram of Skills
  • Sign language: For “I love you” in American sign language, five fingers are usually required, these have to provide flexion/extension DOFs, and the thumb has to be in reposition.
@@
Appendix B Data
Overviews of the analyzed robotic hands and the papers considered for the systematic review are available at https://figshare.com/s/f4fdd9aca1c134f010dd.